# Do different prompting methods yield a common task representation in language models?

### v2

@@ Appendix A Prompt generation prompt templates
{mdframed}
<|start_header_id|>user<|end_header_id|>
Instructions
You are powerful model helping write prompts to help smaller models perform tasks better. Below, you will be given a set of input-output pairs for a particular undescribed task. First, please study the examples to deduce what the task is, and describe your thinking under the header "# Task Deduction". Next, please write 10 prompts that might help a smaller model perform this task. The prompts should be:
  1. 1.
    Short, up to 10 words.
  2. 2.
    Informative about what the task is.
  3. 3.
    Not repetitive with each other.
Please write your prompts under the header "# Task Prompts".
@@
Now, think step by step and follow the instructions above.
<|eot_id|><|start_header_id|>assistant<|end_header_id|>
@@
{mdframed}
<|start_header_id|>user<|end_header_id|>
Instructions
You are powerful model helping write prompts to help smaller models perform tasks better. Below, you will be given a set of input-output pairs for a particular undescribed task. First, please study the examples to deduce what the task is, and describe your thinking under the header "# Task Deduction". Next, please write 10 prompts that might help a smaller model perform this task. The prompts should be:
  1. 1.
    As long as necessary to be helpful for the smaller model.
  2. 2.
    Informative about what the task is.
  3. 3.
    Not repetitive with each other.
  4. 4.
    Not including any examples of the task.
Please write your prompts under the header "# Task Prompts".
@@
Now, think step by step and follow the instructions above.
<|eot_id|><|start_header_id|>assistant<|end_header_id|>
@@ B.1 Additional uninformative instruction details
Ahead of time, we precompute the log-probability of texts from WikiText-103-v1 (Merity et al., , 2016) with each model $f$. We randomly sample entries from the WikiText-103-v1 dataset. For each sampled entry, we extract overlapping prefixes that end with whitespace (excluding the terminal whitespaces), tokenize these prefixes, and keep those with a length of 64 or fewer tokens. We extract prefixes to create strings that might coherently appear at the beginning of a text. We then compute and cache the (log-) probability of each such prefix and its token sequence. We cache approximately $655362^{16}=$ 65536 token sequences with each model (approximately as we stop after the Wikitext entry that brought us over the threshold, but do not discard sequences beyond the $2^{16}$’th one).
To sample an uninformative $\tilde{q}_{m}^{t}$ for some $q_{m}^{t}$, we tokenize $q_{m}^{t}$ to arrive at its length $L(q_{m}^{t})$ and compute its log-probability $\log f(q_{m}^{t})$. We now create a representative sample of $N=100$ cached texts with approximately the same length as $L(q_{m}^{t})$. We begin by only consider texts of with a length of precisely $L(q_{m}^{t})$ tokens. If there are over $N$ of those, we stop; otherwise, for $k\in\mathbb{Z}^{+}$, we also consider texts of length $L(q_{m}^{t})\pm k$, increasing $k$ by one until we have a set of at least $N$ candidates. Once we have attained this set of candidates, we select the $\tilde{q}_{m}^{t}$ with the closest log-probability under $f$ as $q_{m}^{t}$ has, and remove it from the set. As we require five uninformative baselines for each instruction $q_{m}^{t}$ (see Appendix E), we will use the five real texts with the closest log-probability to $q_{m}^{t}$ subject to being within a small number of tokens from $L(q_{m}^{t})$.
@@ Appendix E Full experimental settings
[…]
(Start1=1){easylist}[itemize] @ For each model and each task, we use the $J=5$ instructions with the highest accuracy over the training split. @ We compute the mean activations over 100 total prompts, 20 with each of the 5 best instructions. @ We compute the causal indirect effects over 25 total uninformative prompts, 5 generated for each of the best instructions. @ We batch our results with a batch size that depends on the model and task, but does not exceed 5 for any model or task (see code for batch size computation logic). @ We split each dataset 70% to train and 30% to test. Where we require a validation set, we split it again from the training set. @ We load all models in full precision. @ We use the $|\mathcal{A}|=20$ top heads in all experiments we report. @ We evaluate the FV interventions at every possible depth (that is, after every layer of the model). @@ In all main manuscript figures, we report the accuracy intervening after the $|L/3|$ layer ( (layer 9 for the Llama-3.2-3B models and layer 11 for the Llama-3.1-8B and OLMo-2-1123-7B ones). @@ In appendix figures that report the empirically optimal intervention layer, we compute the accuracy using the layer that would result in the highest mean accuracy for each model, averaging over both the 0-shot and shuffled 10-shot evaluations, separately for demonstrations and instructions. @ When we intervene with two function vectors (either demonstrations and instructions, or twice with one), we sweep over the range $[\left\lfloor L/4\right\rfloor,\left\lceil L/2\right\rceil]$ (as the optimal intervention depths for all models fell in this range. @@ In the main manuscript figure, we report the accuracy intervening with both additively at the $|L/3|$ layer. @@ In appendix figures showcasing the empirically optimal layer(s), we follow the same process described above. @ Error bars we report in all figures are standard errors of the mean, averaged within each model. @ We use the following query template in all of our instruction-based experiments:
{mdframed}
@@
@ And this query template in our demonstration-based ones:
{mdframed}
Q: <$x_{i1}$>
A: <$y_{i1}$>
Q: <$x_{i2}$>
A: <$y_{i2}$>
Q: <$x_{i10}$>
A: <$y_{i10}$>
Q: <$x_{iq}$>
@@ E.1 Experiment compute resources
Before running the main set of experiments, we run a few preliminary steps. We generate instructions (Appendix A) 20 times for each of the $\approx 50$ datasets, generating shorter and longer instructions separately, resulting in approximately 2000 invocations of the stronger model used for instruction generation (in our experiments, Llama-3.1-405B). We also cache the log-probabilities of texts from WikiText-103-v1, which takes an hour or two with each model (Appendix B)
@@
[…]
@@ G.2 Findings 1 and 2 with FV-incongruent evaluations
G.3 Findings 1 additional model results
@@
G.4 Adding an FV twice control
@@ I.1 Finding 4 with empirically optimal layer
Appendix J Finding 4 appendix results
@@
We perform two control experiments to ensure that this observed asymmetry is meaningful (and not an artifact of selecting arbitrary sets of heads). In one, we select the set of heads with the lowest absolute causal scores across both instructions and demonstrations, which we consider to be the heads most unrelated to inducing task representations from either presentation. In another, we select the bottom heads — that is, the ones with the largest negative causal scores for either instructions or prompts. We report both in Section J.1. Performance with the least important heads is indistinguishable from the baselines, and performance with the bottom heads is often below the baselines. The existence of heads with negative causal scores is, itself, curious — these are attention heads whose mean task-conditioned activations lower the probability assigned to the correct token, rather than raise it.

### v3

@@ Appendix A Prompt generation prompt templates
<|start_header_id|>user<|end_header_id|> Instructions You are powerful model helping write prompts to help smaller models perform tasks better. Below, you will be given a set of input-output pairs for a particular undescribed task. First, please study the examples to deduce what the task is, and describe your thinking under the header "# Task Deduction". Next, please write 10 prompts that might help a smaller model perform this task. The prompts should be: 1. Short, up to 10 words. 2. Informative about what the task is. 3. Not repetitive with each other. Please write your prompts under the header "# Task Prompts".
@@
Now, think step by step and follow the instructions above. <|eot_id|><|start_header_id|>assistant<|end_header_id|>
@@
<|start_header_id|>user<|end_header_id|> Instructions You are powerful model helping write prompts to help smaller models perform tasks better. Below, you will be given a set of input-output pairs for a particular undescribed task. First, please study the examples to deduce what the task is, and describe your thinking under the header "# Task Deduction". Next, please write 10 prompts that might help a smaller model perform this task. The prompts should be: 1. As long as necessary to be helpful for the smaller model. 2. Informative about what the task is. 3. Not repetitive with each other. 4. Not including any examples of the task. Please write your prompts under the header "# Task Prompts".
@@
Now, think step by step and follow the instructions above. <|eot_id|><|start_header_id|>assistant<|end_header_id|>
@@ B.1 Additional uninformative instruction details
Ahead of time, we precompute the log-probability of texts from WikiText-103-v1 (Merity et al., , 2016) with each model $f$. We randomly sample entries from the WikiText-103-v1 dataset. For each sampled entry, we extract overlapping prefixes that end with whitespace (excluding the terminal whitespaces), tokenize these prefixes, and keep those with a length of 64 or fewer tokens. We extract prefixes to create strings that might coherently appear at the beginning of a text. We then compute and cache the (log-) probability of each such prefix and its token sequence. We cache approximately $2^{16}=65536$ token sequences with each model (approximately as we stop after the Wikitext entry that brought us over the threshold, but do not discard sequences beyond the $2^{16}$’th one).
To sample an uninformative $\tilde{q}_{m}^{t}$ for some $q_{m}^{t}$, we tokenize $q_{m}^{t}$ to arrive at its length $L(q_{m}^{t})$ and compute its log-probability $\log f(q_{m}^{t})$. We now create a representative sample of $N=100$ cached texts with approximately the same length as $L(q_{m}^{t})$. We begin by only consider texts of with a length of precisely $L(q_{m}^{t})$ tokens. If there are over $N$ of those, we stop; otherwise, for $k\in\mathbb{Z}^{+}$, we also consider texts of length $L(q_{m}^{t})\pm k$, increasing $k$ by one until we have a set of at least $N$ candidates. Once we have attained this set of candidates, we select the $\tilde{q}_{m}^{t}$ with the closest log-probability under $f$ as $q_{m}^{t}$ has, and remove it from the set. As we require five uninformative baselines for each instruction $q_{m}^{t}$ (see Appendix˜E), we will use the five real texts with the closest log-probability to $q_{m}^{t}$ subject to being within a small number of tokens from $L(q_{m}^{t})$.
@@ Appendix E Full experimental settings
• For each model and each task, we use the $J=5$ instructions with the highest accuracy over the training split.
• We compute the mean activations over 100 total prompts, 20 with each of the 5 best instructions.
• We compute the causal indirect effects over 25 total uninformative prompts, 5 generated for each of the best instructions.
• We batch our results with a batch size that depends on the model and task, but does not exceed 5 for any model or task (see code for batch size computation logic).
• We split each dataset 70% to train and 30% to test. Where we require a validation set, we split it again from the training set.
• We load all models in full precision.
• We use the $|\mathcal{A}|=20$ top heads in all experiments we report.
• We evaluate the FV interventions at every possible depth (that is, after every layer of the model).
– In all main manuscript figures, we report the accuracy intervening after the $|L/3|$ layer ( (layer 9 for the Llama-3.2-3B models and layer 11 for the Llama-3.1-8B and OLMo-2-1123-7B ones).
– In appendix figures that report the empirically optimal intervention layer, we compute the accuracy using the layer that would result in the highest mean accuracy for each model, averaging over both the 0-shot and shuffled 10-shot evaluations, separately for demonstrations and instructions.
• When we intervene with two function vectors (either demonstrations and instructions, or twice with one), we sweep over the range $\left\lfloor L/4\right\rfloor,\left\lceil L/2\right\rceil]$ (as the optimal intervention depths for all models fell in this range.
– In the main manuscript figure, we report the accuracy intervening with both additively at the $|L/3|$ layer.
– In appendix figures showcasing the empirically optimal layer(s), we follow the same process described above.
• Error bars we report in all figures are standard errors of the mean, averaged within each model.
• We use the following query template in all of our instruction-based experiments:
@@
• And this query template in our demonstration-based ones:
Q: $x_{i1}$>
A: $y_{i1}$>
Q: $x_{i2}$>
A: $y_{i2}$>
Q: $x_{i10}$>
A: $y_{i10}$>
Q: $x_{iq}$>
@@ E.1 Experiment compute resources
Before running the main set of experiments, we run a few preliminary steps. We generate instructions (Appendix˜A) 20 times for each of the $\approx 50$ datasets, generating shorter and longer instructions separately, resulting in approximately 2000 invocations of the stronger model used for instruction generation (in our experiments, Llama-3.1-405B). We also cache the log-probabilities of texts from WikiText-103-v1, which takes an hour or two with each model (Appendix˜B)
@@
Most of our additional experiments only require these final evaluation jobs:
• Intervening with both function vectors (§3.2; Figure 2B).
• Intervening with the same function vector twice (a control for §3.2; Figure 18).
• Constructing ‘incongruent’ FVs with top heads identified by demonstrations and mean activations from instructions (or vice versa; §3.4; Figure 3G-H).
• Constructing function vectors using the least important overall heads and the bottom heads (a control for §3.4; Figure 30)
• Steering base models with post-trained model instruction FVs (§3.5; Figure 4A).
Many of these we only run for the four models we focus our investigation on (Llama-3.2-3B, Llama-3.2-3B-Instruct, Llama-3.1-8B, Llama-3.1-8B-Instruct). However, all of these run for approximately 50 datasets, using both short and long instruction mean activations. We conservatively estimate these evaluations required another $50$ (tasks) $\times$ 2 (instruction lengths) $\times$ 4 (models) $\times$ 10 (additional experiments) = 4000 experiments.
@@ G.2 Findings 1 and 2 with FV-incongruent evaluations
G.3 Findings 1 and 2 with OLMo models
Figure 13: A replication of Figure 2, with the OLMo models explored in subsection 3.5.
Figure 14: A replication of Figure 2, with the OLMo models explored in subsection 3.5, using the empirically-optimal intervention depth; i.e, Figure 11, but with the OLMo models.
Figure 15: A replication of Figure 2, with the OLMo models explored in subsection 3.5, using the mismatched evaluation version; i.e., Figure 12, but with the OLMo models.
G.4 Findings 1 additional model results
@@
G.5 Adding an FV twice control
@@ I.1 Finding 4 with empirically optimal layer
I.2 Finding 4 with OLMo models
Figure 29: (Left) A replication of the joint intervention results of Figure 3G-H, using the OLMo models explored in subsection 3.5. (Right) Same as the left panel, but using the empirically optimal intervention depths; a version of Figure 28, but with the OLMo models.
Appendix J Finding 4 analyses and controls
@@ I.2 Finding 4 with OLMo models
We perform two control experiments to ensure that this observed asymmetry is meaningful (and not an artifact of selecting arbitrary sets of heads). In one, we select the set of heads with the lowest absolute causal scores across both instructions and demonstrations, which we consider to be the heads most unrelated to inducing task representations from either presentation. In another, we select the bottom heads — that is, the ones with the largest negative causal scores for either instructions or prompts. We report both in Section˜J.1. Performance with the least important heads is indistinguishable from the baselines, and performance with the bottom heads is often below the baselines. The existence of heads with negative causal scores is, itself, curious — these are attention heads whose mean task-conditioned activations lower the probability assigned to the correct token, rather than raise it.